OpenStudy (anonymous):

Which of the following systems of equations has no solution? 1) 8x - 6y = 14; 3y = 4x - 7 2) 6x - 4 = 3y; 6y - 1 = 12x 3) 9x - y = 12; 2y = 18x - 24 4) -5y = x - 8; 2x - 10y = 7

OpenStudy (amistre64):

the ones that are seperate but parallel lines

OpenStudy (amistre64):

look for the 2 that have the same equations; but a different "equals number"

OpenStudy (amistre64):

arrange them so they are both in the same form 8x - 6y = 14 3y = 4x - 7 8x - 6y = 14 -4x + 3y = - 7 [*-2] 8x - 6y = 14 8x -6y = 14 ; these are the same lines

OpenStudy (anonymous):

just use destroying method ...

OpenStudy (amistre64):

6x - 4 = 3y 6y - 1 = 12x 6x -3y = 4 [-2] -12x + 6y = 1 -12x +6y = -8 -12x + 6y = 1 ; that looks like a winner